# Fixed-point Inversion for Text-to-image diffusion models

### v1

@@ Abstract
Text-guided diffusion models offer powerful new ways to generate and manipulate images. Several applications of these models, including image editing interpolation, and semantic augmentation, require diffusion inversion. This is the process of finding a noise seed that can be used to generate a given image. Current techniques for inverting a given image can be slow or inaccurate. The technical challenge for inverting the diffusion process arises from an implicit equation over the latent that cannot be solved in closed form. Previous approaches proposed to solve this issue by approximation or various learning schemes. Here, we formulate the problem as a fixed-point equation problem and solve it using fixed-point iterations, a well-studied approach in numerical analysis. We further identify a source of inconsistency that significantly hurts the inversion of real images encoded to the latent space. We show how to correct it by applying a prompt-aware adjustment of the encoding. Our solution, Fixed-point inversion, is much faster than previous techniques like EDICT [22] and Null-text [13], with similar inversion quality. It can be combined with any pretrained diffusion model and requires no model training, prompt tuning, or additional parameters. In a series of experiments, we find that Fixed-point inversion shows improved results in several downstream tasks: image editing, image interpolation, and generation of rare objects.
@@ 1 Introduction
Text-to-image diffusion models [16, 17, 15, 1] can generate diverse and high-fidelity images based on user-provided text prompts. These models are further used in several important tasks that require inversion, namely, discovering an initial noise (seed) that, when subjected to a forward (denoising) diffusion process along with the prompt, generates the input image. Inversion is used in various tasks including image editing [6], personalizaton [5, 4], seed noise interpolation [18] and generating rare concepts [19].
As inversion became a critical infrastructure block of various tasks, several inversion methods were developed. Primarily, Denoising Diffusion Implicit Models (DDIM) [21] introduced a deterministic sampling technique to map an initial noise seed into a generated image. Conversely, DDIM inversion uses that technique to transform an image into a latent noise representation by approximating the inversion equation. This approximation makes it very fast ($\sim$ 1.5 sec on a single A100 GPU), but it also introduces inconsistencies, (as explained in section 3.2), which causes a noticeable effect in generated images.
Several attempts have already been made to address these inconsistencies in DDIM Inversion. Null-text [13] aims to bring closer the diffusion states of the inversion process and the generation process, by adjusting the unconditional textual embedding used in classifier-free guidance [7]. EDICT [22] uses two coupled noise vectors to invert each other in an alternating fashion, to reach an exact diffusion inversion. Both Null-text and EDICT succeed in providing high-quality reconstruction results, but require a much higher computational cost: storing $3M$ or $16K$ parameters respectively for every inverted image, and running for $20$ or $60$ seconds per inversion, orders of magnitude slower than DDIM inversion.
In this paper, we address the problem of inversion inconsistencies with a fast, simple, yet fidelity-preserving procedure. Building on techniques from the field of numerical solution of equations, we take a simple fixed-point-iteration approach [2]. We name our approach FPI, for Fixed-Point Inversion. FPI enforces the consistent inversion of an image by applying only a few fixed-point iterations at each diffusion step. In particular, we demonstrate how fixed-point iteration theory can be used to solve the implicit functions introduced by DDIM. We show, that in practice, $2$ to $5$ iterations were already sufficient for convergence, yielding significantly more accurate results than DDIM inversion. FPI requires no model training or finetuning, no prompt optimization, or any additional parameters. It can be combined with all pre-trained diffusion models.
We also identify a second source of inconsistency, which significantly hurts inversion of real images. We find that when real images are encoded into the latent space, their encoding may misfit the inversion process since that process was designed to invert generated images for a given prompt. We show that applying a short prompt-aware adjustment of the encoding greatly improves inversion quality of FPI.
Figure 1 illustrates the quality and speed of FPI, for both reconstruction and editing. It inverts real images within 5 seconds, reaching similar or better quality than competing methods while being an order of magnitude faster.
@@
We evaluate FPI extensively. First, we directly assess the quality of inversions found with FPI by measuring reconstruction performance, showing comparable results to [13, 22] but with ×$\times 4$ to ×$\times 12$ speedup gain. We then demonstrate the benefit of FPI in two downstream tasks (1) In Image editing, FPI smoothly changes fine details in the image in a consistent and coherent way, whereas previous methods struggle to do so. (2) In Rare concept generation with [19], and seed interpolation [18], that require diffusion inversion. In both of these tasks, FPI yields more accurate seeds, resulting in superior generated images using [19, 18].
@@ 2 Related work
Text-to-image diffusion models[16, 17, 15, 1] translate random samples (seeds) from a high-dimensional space, guided by a user-supplied text prompt, into corresponding images. DDIM [21], is a widely used deterministic scheduler, that demonstrates the inversion of an image to its latent noise seed. While effective, it was designed for unconditional diffusion models and lacked the capability to handle text prompts or editing. When applied to inversion for text-guided diffusion models, DDIM inversion suffers from inconsistency, particularly when the classifier-free guidance constant is large [13]. This happens due to its reliance on a linear approximation, causing the propagation of errors that result in inaccurate image reconstruction and the loss of content.
Recent studies [13] and [22] Address this limitation that arises from DDIM Inversion. Null-text inversion [13] improves reconstruction accuracy by optimizing the embedding vector of an empty string. This ensures that the diffusion process calculated using DDIM inversion, aligns with the reverse diffusion process.
[12] further incorporates the Null-Text Inversion approach by replacing the null-text embedding with a prompt embedding instead. This enhances convergence time and reconstruction quality but results in inferior image editing quality. In both [13] and [12], the optimized embedding must be stored, resulting in nearly 3 million additional parameters for each image when inverting with StableDiffusion v1.5 for 50 denoising steps.
EDICT [22] introduced invertible neural network layers, specifically Affine Coupling Layers, to compute both backward and forward diffusion paths. However, it comes at the cost of doubling the number of neural function evaluations (NFEs) compared to DDIM inversion, leading to prolonged runtime. Furthermore, additional latent has to be stored for every inverted image ($16K$ additional parameter per image in StableDiffusion [16]).
[23] proposed a novel scheme tailored for EDICT, referred to as bi-directional integration (BDIA). This approach aims to decrease the number of neural function evaluation (NFE) operations in EDICT while preserving accurate diffusion inversion. While it provides examples with strong editing capabilities, it still demands an excessive amount of time (×$\times 10$ than DDIM Inversion).
An alternative method proposed by [8] recommends modifying the noise maps to establish a stronger connection with the provided image, resulting in a non-Gaussian noise distribution. Nevertheless, this approach remains computationally expensive, requiring ×$\times 10$ time than DDIM Inversion. Furthermore, its applicability is restricted solely to editing tasks.
Seed Exploration & Rare concept generation:
Here, we introduce another task that heavily relies on diffusion inversion. Generally, the data used for training diffusion models frequently has an unbalanced distribution of concepts [19]. This makes it hard for diffusion models to generate accurate images of rare concepts, and they often generate distorted images or images of closely related but different concepts (e.g. instead of ”tiger cat”, it generates a ”tiger”, see [19]). To overcome this limitation, SeedSelect [19] employs a small set of real images along with their corresponding seeds for seed optimization. This process iteratively refines a seed until a plausible image is generated by the diffusion model. A follow-up approach, NAO [18], introduced new interpolation paths (NAO-path) and centroids (NAO-centroid) of seeds as initialization for SeedSelect. NAO and SeedSelect both use DDIM Inversions for initial seed evaluation, which highly impacts results.
The current paper offers an alternative for obtaining precise inversion seeds for NAO and SeedSelect, in order to produce better images in terms of appearance and semantic accuracy. It’s important to highlight that existing inversion techniques (Null-Text [13] and EDICT [22]) are not applicable in this context as they necessitate extra parameters for image reconstruction, which impedes the straightforward implementation of interpolation and centroid identification as suggested by NAO.
3 Method
3.1 Preliminaries
@@
We first establish the fundamentals of Diffusion Denoising Probabilistic Models (DDPMs) and Denoising Diffusion Implicit Models (DDIMs). In these models, a forward pass (denoising) is the process that generates an image from a seed noise. A backward pass (inversion) is the process that, given an image, finds its corresponding seed noise.
Diffusion Denoising Probabilistic Models (DDPMs):
@@
DDPMs learn to generate images through a systematic process of iteratively adding Gaussian noise to a data sample until the data distribution is mostly noise. The data distribution is subsequently gradually restored through a reverse diffusion process initiated with a random sample (noise seed) from a Gaussian distribution. In recent, latent diffusion models  [16, 17, 15, 1], noising and denoising are applied to a latent representation of images derived from a variational autoencoder, rather than the images directly. Specifically, the process of mapping a (latent) image to noise is a Markov chain that starts with $z_{0}$, and gradually adds noise to obtain latent variables $z_{1},z_{2},\dots,z_{T}$, following
$q(z_{1},z_{2},\ldots,z_{T}|z_{0})=\Pi^{T}_{t=1}{q(z_{t}|z_{t-1}}).$ (1)
Each step in this process is a Gaussian transition
$q(z_{t}|z_{t-1}):=\mathcal{N}(z_{t},\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}I),$ (2)
parameterized by a schedule $\beta_{0},\beta_{1},\ldots,\beta_{T}\in(0,1)$. $z_{t}$ can be expressed as a linear combination of $z_{0}$ and Gaussian noise $\varepsilon\sim\mathcal{N}(0,I)$:
$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\varepsilon,$ (3)
where $\alpha_{t}=\Pi^{t}_{i=1}(1-\beta_{i})$.
Text-guided diffusion models control the diffusion process with a control signal in the form of a latent representation $p$ of a text prompt (often a CLIP embedding). A denoising network $\epsilon_{\theta}$ aims to minimize the loss:
$\mathcal{L}=\mathbb{E}_{z_{0},\varepsilon\sim\mathcal{N}(0,1),t}\left[||\varepsilon-\epsilon_{\theta}(z_{t},t,p)||_{2}^{2}\right].$ (4)
The role of the denoising network $\epsilon_{\theta}$ is to accurately eliminate the added noise $\varepsilon$ from the latent code $z$ at every time step $t$, based on the input of the noisy latent code $z_{t}$ and the encoding of a conditioning text $p$.
Denoising Diffusion Implicit Models (DDIM).
@@
Sampling from diffusion models can be viewed alternatively as solving the corresponding diffusion Ordinary Differential Equations (ODEs) [10]. DDIM [21] scheduler, a popular deterministic scheduler, proposed denoising a latent noise vector in the following way:
@@
The noise term $\varepsilon$ that was added in Eq. (3) is replaced here with the output of the network $\epsilon_{\theta}(z_{t},t,p)$ that was trained to predict it.
Figure 1: Fixed-Point Inversion iterates over an implicit function (Eq. 8), at every time step in the backward path. It starts with $z^{0}_{t}=z_{t-1}$ and quickly converges to $z_{t}$. Each box denotes one inversion step; black circles correspond to intermediate latents in the denoising process; green circles correspond to intermediate fixed point iterations.
DDIM inversion:
We now focus on inversion in the latent representation. Given an image representation $z_{0}$ and its corresponding text prompt $p$, we seek a noise seed $z_{T}$ that, when denoised, reconstructs the latent $z_{0}$. Several approaches were proposed for this task, and we focus on DDIM inversion. In this technique, Eq. (5) is rewritten as:
$z_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t-1}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}(z_{t},t,p).$ (6)
This is an implicit equation in $z_{t}$ that can be approximated by replacing $z_{t}$ with $z_{t-1}$
$\approx\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t-1}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}({\color[rgb]{1,0,0}z_{t-1}},t,p).$ (7)
@@
By applying Eq. (7) repeatedly for every denoising step $t$, one can invert an image latent $z_{0}$ to a latent $z_{T}$ in the seed space. DDIM inversion is fast, but the approximation of Eq.( 7) inherently introduces errors at each time step. As these errors accumulate, they cause the whole diffusion process to become inconsistent in the forward and the backward processes, leading to poor image reconstruction and editing [13, 22].
3.2 Fixed-point Inversion
We propose to replace the approximation of the inversion process of Eq. (7), by solving the implicit equation Eq. (6) directly. For that purpose, we propose using fixed point iterations, a common technique in the optimization of implicit equations [2]. Specifically, using upper-script $i$ to denote iterations, we have:
$z_{t}^{0}$ $=$ $z_{t-1}$ (8)
$z_{t}^{i+1}$ $=$ $f(z_{t}^{i}),$
where
$f(z_{t})=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t-1}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}(z_{t},t,p).$ (9)
That is, given a $z_{t-1}$ we perform fixed-point iterations to minimize the residual error of the implicit equation.
Figure 1 illustrates this process. The output of our method is a single latent $z$, similar to the DDIM Inversion’s output. Note that DDIM becomes a unique case of our scheme where the initial estimation $z_{t}^{0}$ is considered as the final output. As we show below, it leads to an improved reconstruction of the image in terms of PSNR with minimal run time overhead.
An advantage of our method is that we only modify the inversion process without altering the forward process of the diffusion model. This makes our method very easy to integrate into all diffusion models. Furthermore, this ensures that the reconstruction (forward) process remains as fast as the DDIM basic process, as it has not been altered.
More sophisticated methods are known for solving implicit functions like Newton-Raphson and Conjugate Gradient [14]. We found these methods to be impractical in this case due to the high dimensionality of the seed space or because they needed much more time to converge. Consequently, we opted for a simpler approach, which already achieves accurate solutions within a few iterations. Further details can be found in the supplementary material.
Figure 2: Fixed-Point Iterations: We measure the effect of the number of fixed-point iterations. (Left) The log loss of the error $f(x)-x$ for 50 images (plot for each image). It shows that FP-inversion converges for all images we tested. (Right) PSNR of reconstructed images, vs the maximum number of FP iterations. The dashed line represents the upper bound induced by Stable Diffusion VAE. It shows that FPI converges in only 2-3 iterations, reaching a PSNR closely approaching the upper limit established by the diffusion model VAE.
Convergence.
[…]
In our case, it is not theoretically guaranteed that the function $f$ is contractive everywhere. Luckily, the mapping may still be contractive in our region of interest. Indeed, we find that in practice, iterating Eq. (6) reduces the difference $\Delta z=z_{t}^{i}-f(z_{t}^{i})$ (the equation residual) along the iterations. Figure 2(left) shows FPI loss (residuals) curves as a function of the number of iterations on a single diffusion step. Figure 2(right) depicts the effect of the number of running iterations on image reconstruction, in terms of PSNR. The figures illustrate that FPI converges in only 2-3 iterations, achieving a PSNR close to the upper bound set by the diffusion model VAE.
3.3 Fixed-point iterations for consistent inversion
We now focus on an important subtlety that greatly impacts inversion quality. Our inversion process is designed to agree with a generation process for a prompt $p$. However, inversion is typically applied to photos and images that are not generated by the model. Such an image $x$ is mapped into the latent $z_{0}$ in a way that is unaware of the prompt because the VAE encoder is prompt-agnostic. This means that the typical input to our inversion process may be inconsistent with the prompt we use.

### v2

# Regularized Newton Raphson Inversion for Text-to-Image Diffusion Models

@@ Abstract
Diffusion inversion is the problem of taking an image and a text prompt that describes it and finding a noise latent that would generate the image. Most current inversion techniques operate by approximately solving an implicit equation and may converge slowly or yield poor reconstructed images. Here, we formulate the problem as finding the roots of an implicit equation and design a method to solve it efficiently. Our solution is based on Newton-Raphson (NR), a well-known technique in numerical analysis. A naive application of NR may be computationally infeasible and tends to converge to incorrect solutions. We describe an efficient regularized formulation that converges quickly to a solution that provides high-quality reconstructions. We also identify a source of inconsistency stemming from prompt conditioning during the inversion process, which significantly degrades the inversion quality. To address this, we introduce a prompt-aware adjustment of the encoding, effectively correcting this issue. Our solution, Regularized Newton-Raphson Inversion, inverts an image within 0.5 sec for latent consistency models, opening the door for interactive image editing. We further demonstrate improved results in image interpolation and generation of rare objects.
Figure 1: Image editing using our approach for inversion demonstrates significant speed-up and improved quality compared to previous state-of-the-art methods. Results are shown for both Latent Diffusion models and fast Latent consistency models.
@@ 1 Introduction
Text-to-image diffusion models [26, 27, 25, 1] can generate diverse and high-fidelity images based on user-provided text prompts. These models are further used in several important tasks that require inversion, namely, discovering an initial noise (seed) that, when subjected to a backward (denoising) diffusion process along with the prompt, generates the input image. Inversion is used in various tasks including image editing [11], personalizaton [9, 8], seed noise interpolation [28], for semantic augmentation and generating rare concepts [29].
As inversion became a critical building block in various tasks, several inversion methods have been suggested. Denoising Diffusion Implicit Models (DDIM) [32] introduced a deterministic and fast sampling technique for image generation with diffusion models. However, DDIM inversion transforms an image back into a latent noise representation by approximating the inversion equation. Although this approximation makes it very fast, it also introduces an approximation error (as explained in section 3), causing noticeable distortion artifacts in inverted images. This is particularly noticeable in consistency models [19, 30], with large gaps between the diffusion time-steps, and where inference is achieved with only 2 – 4 DDIM steps. Several attempts have been made to address the inconsistencies in DDIM Inversion  [22, 33, 23], which lead to poor reconstruction. AIDI [23] solved the DDIM-inversion implicit equation using fixed-point iterations  [23], a numerical scheme with linear convergence rate, while [12] addresses the minimization of the residual error in that equation by gradient descent. Despite these methods showing improvements over previous approaches, their reconstruction and editing performance remain poor and often require extensive amounts of time.
In this paper, we frame the diffusion inversion problem as finding specific roots of an implicit function and propose a solution based on the Newton-Raphson (NR) numerical scheme [16]. Widely used in numerical analysis for its robustness and speed, NR (theoretical) quadratic convergence rate [3, 2, 6], enables very fast inversion. However, we find that a naive application of standard NR often fails to find the correct root, leading to significant distortions in the reconstructed images. We show that this issue can be resolved by regularizing the objective with a prior, derived from the training process of diffusion models. We name our approach RNRI, for Regularized Newton Raphson Inversion. RNRI converges to a consistent inversion of an image in a small number of steps at each diffusion stage. In practice, 1-2 iterations are sufficient for convergence that yields significantly more accurate results than other inversion methods. RNRI requires no model training or finetuning, no prompt optimization, or any additional parameters. It can be combined with all pre-trained diffusion models, and we demonstrate its benefits to inversion of latent diffusion models (LDM) [26] and of latent consistency models (LCM) [30]. Figure 1 demonstrates the quality and speed of RNRI for editing, compared to a SoTA inversion method. Using latent diffusion models that require 50 DDIM steps, our approach can edit real images within 1.9 seconds. For latent consistency models requiring just 4 DDIM steps, the process converges in only 0.5 seconds. To the best of our knowledge, this speed allows users to edit images on-the-fly, using text-to-image models, for the first time.
In practice, high-quality editing using LDM or LCM requires applying classifier-free guidance with a large guidance scale. Similar to [22, 24, 13], we also identify that inversion with high guidance scale significantly degrades the results. When images are encoded into the latent space their encoding often miss-aligns with the prompt. We demonstrate that applying a brief adjustment to the encoding substantially enhances the inversion quality of all iterative methods.
@@
We evaluate RNRI extensively. First, we directly assess the quality of inversions found with RNRI by measuring reconstruction errors, showing comparable results to [22, 33] but with ×$\times 4$ to ×$\times 12$ speedup gain. We then demonstrate the benefit of RNRI in two downstream tasks (1) In Image editing, RNRI smoothly changes fine details in the image in a consistent and coherent way, whereas previous methods struggle to do so. With RNRI inversion, we achieve the fastest editing times, where in the case of latent consistency models a single edit takes less than 0.5 seconds. This capability significantly enhances real-time editing possibilities. (2) In Rare concept generation with [29], and seed interpolation [28], that require diffusion inversion. In both of these tasks, RNRI yields more accurate seeds, resulting in superior generated images, both qualitatively and quantitatively, using the methods in [29, 28].
@@ 2 Related work
Text-to-image diffusion models[26, 27, 25, 1] translate random samples (seeds) from a high-dimensional space, guided by a user-supplied text prompt, into corresponding images. DDIM [32] is a widely used deterministic scheduler, that demonstrates the inversion of an image to its latent noise seed. When applied to inversion of text-guided diffusion models, DDIM inversion suffers from low reconstruction accuracy that is reflected in further tasks, particularly when the classifier-free guidance constant is large [22]. This happens because it relies on a linear approximation, causing a propagation of errors that result in inaccurate image reconstruction and the loss of content. Recent studies [22, 33, 23, 13] address this limitation. Null-text inversion [22] optimizes the embedding vector of an empty string. This ensures that the diffusion process calculated using DDIM inversion, aligns with the reverse diffusion process. [21] replace the null-text embedding with a prompt embedding instead. This enhances convergence time and reconstruction quality but results in inferior image editing quality. In both [22] and [21], the optimized embedding must be stored, resulting in nearly 3 million additional parameters for each image (using 50 denoising steps of StableDiffusion  [26]).
EDICT [33] introduced invertible neural network layers, specifically Affine Coupling Layers, to compute both backward and forward diffusion paths. However effective, it comes at the cost of prolonging inversion time. BDIA [34] introduced a novel integration approximation designed for EDICT, enhancing its computational efficiency while maintaining accurate diffusion inversion. Nevertheless, it still requires significantly more time (10 times longer than DDIM Inversion). AIDI [23] uses an accelerated fixed-point iteration technique at each inversion step to address the implicit function posed by DDIM equations. ExactDPM [13] utilizes gradient-based methods to achieve an effective inversion. ReNoise [10] further extends the work of [23] for Latent consistency models.
Alternative approaches proposed in [15, 4] suggest inverting with a DDPM scheduler instead of DDIM. They begin by constructing auxiliary images $x_{1},...,x_{T}$ and then extracting noise maps $z_{1},...,z_{T}$ to achieve error-free image reconstruction. Despite their effectiveness, the stochastic nature of these methods presents several challenges. Firstly, it necessitates storing $T+1$ latents for every inverted image, resulting in an additional $16K\times T$ parameters. Secondly, the method is solely applicable to reconstruction and editing tasks. Lastly, DDPM inversion typically yields lower-quality results compared to DDIM inversion  [14]. Since our focus is on DDIM deterministic inversion, these papers fall outside the scope of this paper.
3 Preliminaries
@@
We first establish the fundamentals of Denoising Diffusion Implicit Models (DDIMs). In this model, a backward pass (denoising) is the process that generates an image from a seed noise. A forward pass is the process of adding noise gradually to an image until it becomes pure Gaussian noise. Inversion is similar to the forward pass but the goal is to end with a specific Gaussian noise that would generate the image if denoised.
@@
Forward Pass in Diffusion Models. Latent diffusion and consistency models learn to generate images through a systematic process of iteratively adding Gaussian noise to a latent data sample until the data distribution is mostly noise. The data distribution is subsequently gradually restored through a reverse diffusion process initiated with a random sample (noise seed) from a Gaussian distribution. In more detail, the process of mapping a (latent) image to noise is a Markov chain that starts with $z_{0}$, and gradually adds noise to obtain latent variables $z_{1},z_{2},\dots,z_{T}$, following $q(z_{1},z_{2},\ldots,z_{T}|z_{0})=\Pi^{T}_{t=1}{q(z_{t}|z_{t-1}})$, where $\forall t:z_{t}\in\mathbb{R}^{d}$ with $d$ denoting the dimension of the space. Each step in this process is a Gaussian transition
$q(z_{t}|z_{t-1})\sim\mathcal{N}(z_{t},\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}I),$ (1)
parameterized by a schedule $\beta_{0},\beta_{1},\ldots,\beta_{T}\in(0,1)$.
@@
Denoising Diffusion Implicit Models (DDIM). Sampling from diffusion models can be viewed alternatively as solving the corresponding diffusion Ordinary Differential Equations (ODEs) [18]. DDIM [32] scheduler, a popular deterministic scheduler, proposed denoising a latent noise vector in the following way:
@@
$\epsilon_{\theta}(z_{t},t,p)$ is the output of a network that was trained to predict the noise to be removed.
DDIM inversion. We now focus on inversion in the latent representation. Given an image representation $z_{0}$ and its corresponding text prompt $p$, we seek a noise seed $z_{T}$ that, when denoised, reconstructs the latent $z_{0}$. Several approaches were proposed for this task, and we focus on DDIM inversion. In this technique, Eq. (2) is rewritten as:
$z_{t}$ $=$ $f(z_{t})$ (3)
$f(z_{t})$ ::=:= $\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}(z_{t},t,p).$
DDIM inversion approximates this implicit equation in $z_{t}$ by replacing $z_{t}$ with $z_{t-1}$
$\approx\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t}}\cdot\Delta\psi(\alpha_{t})\cdot\epsilon_{\theta}({\color[rgb]{1,0,0}z_{t-1}},t,p).$ (4)
@@
By applying Eq. (4) repeatedly for every denoising step $t$, one can invert an image latent $z_{0}$ to a latent $z_{T}$ in the seed space. DDIM inversion is fast, but the approximation of Eq.( 4) inherently introduces errors at each time step. As these errors accumulate, they cause the whole diffusion process to become inconsistent in the forward and the backward processes, leading to poor image reconstruction and editing [22, 33, 23]. This is particularly noticeable in consistency models with a small number of DDIM steps (typically 2-4 steps), where there’s a significant gap between $z_{t}$ and $z_{t-1}$, see Figure 5(b).
Iterative inversion optimization methods: Several papers proposed to improve the approximation using iterative methods [23, 13]. AIDI [23] proposed to directly solve Eq.(3) using fixed-point iterations [5], a widely-used method in numerical analysis for solving implicit functions. They solve $z=f(z)$ using fixed-point iterations. In a related way, [13] solves a more precise inversion equation, obtained by employing higher-order terms, using gradient descent.
Figure 2: Newton-Raphson Inversion iterates over an implicit function (Eq. 8), at every time step in the inversion path. It starts with $z^{0}_{t}=z_{t-1}$ and quickly converges (within 2 iterations) to $z_{t}$. Each box denotes one inversion step; black circles correspond to intermediate latents in the denoising process; green circles correspond to intermediate Newton-Raphson iterations.
4 Our method: Regularized Newton Raphson Inversion
DDIM inversion above often yields poor reconstruction accuracy. Proposed improvements described above have a linear convergence rate and may take many seconds to compute. In this paper we describe a faster and more robust alternative based on the well-known Newton-Raphson method (NR) [16]. Newton-Raphson is a method for iteratively finding the roots of a system of equations.
A naive Newton-Raphson approach. Consider first a naive way to apply NR to the inversion problem. We can define the following vector residual function,
$r(z_{t}):=z_{t}-f(z_{t})$ (5)
where $r:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, and then seek its zero-crossing roots. That is, find those roots $z_{t}$ for which $r(z_{t})=\mathbf{0}$, meaning that the two vectors $z_{t}$, $f(z_{t})$ are identical. For this problem, NR operates in the following way [6]. Given an initial guess $z_{0}$, It iterates $z^{k+1}_{t}=z^{k}_{t}-\left(J(z^{k}_{t})\right)^{-1}\cdot{r}(z^{k}_{t})$, where $(J(z_{t}^{k}))^{-1}$ presents the inverse of a Jacobian matrix $J\in\mathbb{R}^{d\times d}$ (all derivatives are w.r.t to $z$). In our case, applying this naive scheme is impractical, because the dimension of $z$ is high ($d\approx 16K$ in StableDiffusion [26]), making it too expensive to compute the Jacobian in terms of time- and memory cost, and inverting it becomes practically infeasible.
To address this computational limitation, we propose to apply NR to a multi-variable, scalar function instead. Specifically, apply a norm over $r(z_{t})$, $\hat{r}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}_{0}$:
$\hat{r}(z_{t})$ ::=:= $||z_{t}-f(z_{t})||$ (6)
and seek for solutions that satisfy $\hat{r}(z_{t})=0$. Here, |||||$\cdot$||| | ⋅ | | denotes a norm; we used $L_{1}$ to simply sum over all absolute values of $\hat{r}$. This reduction transforms the Jacobian matrix into a vector, making it easy and fast for computation (see derivation in Appendix A). Under some conditions, and particularly when there is only a single root in the $\epsilon$-neighborhood of the initial guess, the Newton-Raphson method can be proved to have a quadratic convergence rate [16]. In fact, $\hat{r}$ introduces an underdetermined equation that may have multiple roots, and in practice, we find that solving Eq. 6 for $r(z_{t})=0$ often converges to a solution that is out-of-distribution for the diffusion model. This inversion solution leads to bad reconstruction. We address this issue in the next section.
Figure 3: (a) Convergence rate. Comparison of iterative methods in an image inversion-reconstruction task over the COCO validation set. The mean PSNR of reconstructed images is plotted against the number of iterations. The dashed line represents the upper bound on reconstruction quality determined by the VAE in Stable Diffusion. Mean convergence time (in seconds) is denoted for each method. Our RNRI achieves a PSNR close to the upper limit and converges within only 1-2 iterations. (b) Prior effect on convergence. Incorporating our prior not only aids in finding the correct solution but also accelerates convergence.
4.1 Regularized Newton Raphson
To address the undetermined nature of our equation we suggest adding a regularization term, that can be viewed also as a soft constraint, to the Newton-Raphson objective. More precisely, since each step in the diffusion process follows a Gaussian distribution $q(z_{t}|z_{t-1})$ (Sec 3, Eq. 1), we add a prior over the values of $z_{t}$, by adding the negative log-likelihood as a regularizing penalty term to the objective. The objective is thus:
$\mathcal{F}(z_{t}):=||z_{t}-f(z_{t})||_{1}-\lambda\sum_{j=1}^{d}\log q_{j}(z_{t}|z_{t-1})$ (7)
Here, $\lambda>0$ is a hyperparameter weighting factor for the regularization, and $q_{j}$ is the $j$’th component of $q(z_{t}|z_{t-1})$ (refer to Appendix D for more details). We now seek a solution that satisfies $\mathcal{F}(z_{t})=0$. Note that this equation is strictly satisfied iff the residual $\hat{r}(z_{t})=0$ and the probabilities $q_{j}(z_{t}|z_{t-1})=1$, for all $j$.
The Newton Raphson iteration scheme for finding roots of our scalar function in Eq. (7) is given by (see derivation in Appendix A):
$z_{t}^{0}$ $=z_{t-1}$
$z_{t}^{k+1}(i)$ $=z_{t}^{k}(i)-\frac{1}{d}~{}\frac{{\mathcal{F}}(z_{t}^{k})}{g_{i}(z_{t}^{k})+\eta}.$ (8)
Here $g_{i}:=\frac{\partial\mathcal{F}(z_{t})}{\partial z_{t}(i)}$ is the partial derivative of $\mathcal{F}$ with respect to the variable (descriptor) $z_{t}(i)$. $\eta$ is a small constant added for numerical stability and $i\in[0,1,...,d]$ indicates $z$’s components. $g$ can be computed efficiently using automatic differentiation engines. We initialize the process with $z$ from the previous diffusion timestep. Figure 2 illustrates this process. Note that while the solution of Eq. 7 matches the minimizer of $\mathcal{F}$, we employ the NR scheme to solve the equation. We call our approach Regualarized Newton Raphson Inversion (RNRI). Figure 3 (left) depicts the effect of the number of running iterations on image reconstruction, in terms of PSNR, compared to other iterative methods. The figures illustrate that RNRI converges in only 1-2 iterations, achieving a PSNR close to the upper bound set by the diffusion model VAE. Figure 3 (right) shows the impact of prior on the residual. We depict residual curves for 50 random COCO images. Regularization yields superior outcomes, as evidenced by lower residuals.
4.2 Prompt-aware adjustment for consistent inversion